find the exact value of the indicated trigonometric function of omega, if omega is an angle in standard position whose terminal side contains the given point.
(12,16); find sin omega

Answer by lwsshak3(11628)

find the exact value of the indicated trigonometric function of omega, if omega is an angle in standard position whose terminal side contains the given point.
(12,16); find sin omega
***
You are working with a reference right triangle in quadrant I
hypotenuse=√(12^2+16^2)=√400=20
sin omega=16/20=4/5
cos omega=12/20=3/5
tan omega=sin/cos=4/3